# Simultaneous Motion And Noise Estimation with Event Cameras

### v1

@@ 6.1 Video
dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ $5$ Hz Noise CMax [17] 20.000 709 6820.0007096820.000\,709\,6820.000 709 68 1.258 796 0971.2587960971.258\,796\,0971.258 796 097 124.641 377124.641377124.641\,377124.641 377 1.128 529 5931.1285295931.128\,529\,5931.128 529 593 – w/ Init. 8.274 876 2458.2748762458.274\,876\,2458.274 876 245 1.273 697 0451.2736970451.273\,697\,0451.273 697 045 20.659 240 3320.6592403320.659\,240\,3320.659 240 33 1.214 159 4591.2141594591.214\,159\,4591.214 159 459 Downsampling $90$% 8.807 703 958.807703958.807\,703\,958.807 703 95 1.272 907 8221.2729078221.272\,907\,8221.272 907 822 124.868 639 7124.8686397124.868\,639\,7124.868 639 7 1.107 758 3951.1077583951.107\,758\,3951.107 758 395 – w/ Init. 8.226 219 4448.2262194448.226\,219\,4448.226 219 444 1.273 725 3071.2737253071.273\,725\,3071.273 725 307 20.618 615 8620.6186158620.618\,615\,8620.618 615 86 1.214 188 4551.2141884551.214\,188\,4551.214 188 455 Downsampling $80$% 9.964 567 2999.9645672999.964\,567\,2999.964 567 299 1.270 984 5391.2709845391.270\,984\,5391.270 984 539 161.810 797 8161.8107978161.810\,797\,8161.810 797 8 1.060 671 $1$.0606711111.060\,671\,$1$.060 671 111 – w/ Init. 8.243 565 3688.2435653688.243\,565\,3688.243 565 368 1.273 741 8421.2737418421.273\,741\,8421.273 741 842 20.797 769 5620.7977695620.797\,769\,5620.797 769 56 1.214 169 4341.2141694341.214\,169\,4341.214 169 434 Downsampling $70$% 14.398 878 9114.3988789114.398\,878\,9114.398 878 91 1.262 100 3441.2621003441.262\,100\,3441.262 100 344 184.511 383 4184.5113834184.511\,383\,4184.511 383 4 1.030 485 5451.0304855451.030\,485\,5451.030 485 545 – w/ Init. 8.231 444 7738.2314447738.231\,444\,7738.231 444 773 1.273 783 9431.2737839431.273\,783\,9431.273 783 943 23.678 913 0923.6789130923.678\,913\,0923.678 913 09 1.210 643 261.210643261.210\,643\,261.210 643 26 Ours $90$% 8.521 574 0738.5215740738.521\,574\,0738.521 574 073 1.273 446 2961.2734462961.273\,446\,2961.273 446 296 97.584 574 8497.5845748497.584\,574\,8497.584 574 84 1.140 880 661.140880661.140\,880\,661.140 880 66 – w/ Init. 8.188 607 2268.1886072268.188\,607\,2268.188 607 226 1.273 725 2991.2737252991.273\,725\,2991.273 725 299 20.604 334 6220.6043346220.604\,334\,6220.604 334 62 1.214 179 0181.2141790181.214\,179\,0181.214 179 018 Ours $80$% 8.511 006 6718.5110066718.511\,006\,6718.511 006 671 1.273 293 7311.2732937311.273\,293\,7311.273 293 731 131.355 614 8131.3556148131.355\,614\,8131.355 614 8 1.102 904 5081.1029045081.102\,904\,5081.102 904 508 – w/ Init. 8.170 039 698.170039698.170\,039\,698.170 039 69 1.273 790 7891.2737907891.273\,790\,7891.273 790 789 20.861 568 0320.8615680320.861\,568\,0320.861 568 03 1.214 094 851.214094851.214\,094\,851.214 094 85 Ours $70$% 9.180 407 8239.1804078239.180\,407\,8239.180 407 823 1.272 074 7881.2720747881.272\,074\,7881.272 074 788 163.676 240 5163.6762405163.676\,240\,5163.676 240 5 1.061 940 9851.0619409851.061\,940\,9851.061 940 985 – w/ Init. 8.086 262 248.086262248.086\,262\,248.086 262 24 1.273 578 8461.2735788461.273\,578\,8461.273 578 846 21.151 118 9621.1511189621.151\,118\,9621.151 118 96 1.213 969 0641.2139690641.213\,969\,0641.213 969 064 BAF 19.674 893 0519.6748930519.674\,893\,0519.674 893 05 1.259 772 491.259772491.259\,772\,491.259 772 49 125.028 145 1125.0281451125.028\,145\,1125.028 145 1 1.127 308 9021.1273089021.127\,308\,9021.127 308 902 – w/ Init. 8.253 078 2028.2530782028.253\,078\,2028.253 078 202 1.273 636 991.273636991.273\,636\,991.273 636 99 19.550 436 619.550436619.550\,436\,619.550 436 6 1.213 505 481.213505481.213\,505\,481.213 505 48 $1$ Hz CMax [17] 19.395 190 2319.3951902319.395\,190\,2319.395 190 23 1.275 866 1481.2758661481.275\,866\,1481.275 866 148 117.439 798 6117.4397986117.439\,798\,6117.439 798 6 1.143 852 7451.1438527451.143\,852\,7451.143 852 745 – w/ Init. 8.254 263 2638.2542632638.254\,263\,2638.254 263 263 1.289 966 6551.2899666551.289\,966\,6551.289 966 655 20.627 507 7120.6275077120.627\,507\,7120.627 507 71 1.223 032 6881.2230326881.223\,032\,6881.223 032 688 Downsampling $90$% 8.676 346 0658.6763460658.676\,346\,0658.676 346 065 1.289 459 4391.2894594391.289\,459\,4391.289 459 439 110.568 366 3110.5683663110.568\,366\,3110.568 366 3 1.130 495 7731.1304957731.130\,495\,7731.130 495 773 – w/ Init. 8.184 187 8618.1841878618.184\,187\,8618.184 187 861 1.289 611 9121.2896119121.289\,611\,9121.289 611 912 20.620 358 2120.6203582120.620\,358\,2120.620 358 21 1.223 050 3321.2230503321.223\,050\,3321.223 050 332 Ours $90$% 8.505 807 8288.5058078288.505\,807\,8288.505 807 828 1.289 787 0891.2897870891.289\,787\,0891.289 787 089 87.774 541 8687.7745418687.774\,541\,8687.774 541 86 1.158 678 1341.1586781341.158\,678\,1341.158 678 134 – w/ Init. 8.177 036 3548.1770363548.177\,036\,3548.177 036 354 1.290 007 1851.2900071851.290\,007\,1851.290 007 185 20.569 191 6120.5691916120.569\,191\,6120.569 191 61 1.223 052 861.223052861.223\,052\,861.223 052 86 BAF 19.568 905 3719.5689053719.568\,905\,3719.568 905 37 1.275 738 1221.2757381221.275\,738\,1221.275 738 122 117.554 331 4117.5543314117.554\,331\,4117.554 331 4 1.143 238 0781.1432380781.143\,238\,0781.143 238 078 – w/ Init. 8.189 429 3938.1894293938.189\,429\,3938.189 429 393 1.289 532 0351.2895320351.289\,532\,0351.289 532 035 20.713 437 520.713437520.713\,437\,520.713 437 5 1.223 006 3741.2230063741.223\,006\,3741.223 006 374
@@ 6.2 Denoising Convergence During Optimization
To further validate the proposed joint estimation approach, we analyze the convergence during the joint estimation using the ECD dataset in Fig. 11 (see also results in Fig. 5). At the first iteration (i.e., initialization), signal and noise events are classified randomly. As the optimization proceeds, signal events evolve to keep edge structures in the scene, while noise events evolve to drop such edge structures (see the second and third rows). Also, the IWEs converge to sharp edges with correct motion parameters. This example further confirms the efficacy of the joint estimation.
The optimization process on an HD-resolution ($1280\times 720$ px) real-world dataset, TUM-VIE [32], is shown in Fig. 12. The intermediate progress (d)–(e) demonstrates how motion estimation converges and the denoising performance improves, simultaneously.
@@ 6.4 Full Results on Angular Velocity Estimation
While the quantitative evaluation on angular velocity estimation is summarized in Fig. 6, here, we report the detailed results with different target ratio parameters, also compared with other baselines such as BA Filter [9]. The original CMax degrades due to noise, as reported in previous work (e.g., [3]). The S/N target ratio $\tau$ affects the accuracy: when it is close to the actual value of noise injection, the result of the proposed method becomes better. The amount of artificial noise injected is around $15$ for $5$ Hz and $3$ for $1$ Hz conditions. Although we cannot know the “true” noise level due to the original noise in the ECD sequences, our method constantly produces better accuracy and FWL than other baselines. Please refer to Sec. 4.3 for more discussions about dependency on initialization and comparison with other baselines. The AUCs for the conditions that we test ($\tau=\{0.9,\ldots,0.7\}$) are $0.70$ (“Ours”) and $0.67$ (“Downsampling”).
@@ 6.5 Quantitative Evaluation of Intensity Reconstruction
Nonetheless, we report non-reference image quality indices for different S/N ratios ($\tau$). Figure 13 reports the scores of Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [37] and Naturalness Image Quality Evaluator (NIQE) [38], using Bicycle-ND64-2 sequence (same as Fig. 8). These scores indicate the perceptual quality of images, and smaller is better. Although BRISQUE monotonically increases as the target ratio decreases (i.e., more events are removed), NIQE scores the lowest at $\tau=0.9$, indicating the best quality of the reconstructed image. Although the results potentially suggest that it could estimate the “true” noise ratio in the data using the non-reference indices, which is useful for image reconstruction applications, we leave further evaluation and discussion as future work.
Figure 13: Results of the non-reference image quality indices for image reconstruction. See Fig. 8 for the details.
7 Limitations
Our method expects as parameter a target noise level, but the intrinsic (“true”) noise level depends on the camera, scene, etc. and is therefore unknown. More research in this direction is needed for automatic parameter tuning.
Also, notice that there might be a potential limitation of the denoising metric ESR, since it depends on the motion parameter (see Suppl. Mat. of [28]). Although our denoising efficacy is confirmed from multiple perspectives, we believe that advancing denoising metrics for real-world data (where it gets interesting, as there are no GT labels) is a central direction for future work in event-based vision.

### v2

@@ 7.1 Video
dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ 555 Hz Noise CMax [17] 20.000 709 6820.000\,709\,6820.000 709 68 1.258 796 0971.258\,796\,0971.258 796 097 124.641 377124.641\,377124.641 377 1.128 529 5931.128\,529\,5931.128 529 593 – w/ Init. 8.274 876 2458.274\,876\,2458.274 876 245 1.273 697 0451.273\,697\,0451.273 697 045 20.659 240 3320.659\,240\,3320.659 240 33 1.214 159 4591.214\,159\,4591.214 159 459 Downsampling 909090% 8.807 703 958.807\,703\,958.807 703 95 1.272 907 8221.272\,907\,8221.272 907 822 124.868 639 7124.868\,639\,7124.868 639 7 1.107 758 3951.107\,758\,3951.107 758 395 – w/ Init. 8.226 219 4448.226\,219\,4448.226 219 444 1.273 725 3071.273\,725\,3071.273 725 307 20.618 615 8620.618\,615\,8620.618 615 86 1.214 188 4551.214\,188\,4551.214 188 455 Downsampling 808080% 9.964 567 2999.964\,567\,2999.964 567 299 1.270 984 5391.270\,984\,5391.270 984 539 161.810 797 8161.810\,797\,8161.810 797 8 1.060 671 $1$.060\,671\,$1$.060 671 111 – w/ Init. 8.243 565 3688.243\,565\,3688.243 565 368 1.273 741 8421.273\,741\,8421.273 741 842 20.797 769 5620.797\,769\,5620.797 769 56 1.214 169 4341.214\,169\,4341.214 169 434 Downsampling 707070% 14.398 878 9114.398\,878\,9114.398 878 91 1.262 100 3441.262\,100\,3441.262 100 344 184.511 383 4184.511\,383\,4184.511 383 4 1.030 485 5451.030\,485\,5451.030 485 545 – w/ Init. 8.231 444 7738.231\,444\,7738.231 444 773 1.273 783 9431.273\,783\,9431.273 783 943 23.678 913 0923.678\,913\,0923.678 913 09 1.210 643 261.210\,643\,261.210 643 26 Ours 909090% 8.521 574 0738.521\,574\,0738.521 574 073 1.273 446 2961.273\,446\,2961.273 446 296 97.584 574 8497.584\,574\,8497.584 574 84 1.140 880 661.140\,880\,661.140 880 66 – w/ Init. 8.188 607 2268.188\,607\,2268.188 607 226 1.273 725 2991.273\,725\,2991.273 725 299 20.604 334 6220.604\,334\,6220.604 334 62 1.214 179 0181.214\,179\,0181.214 179 018 Ours 808080% 8.511 006 6718.511\,006\,6718.511 006 671 1.273 293 7311.273\,293\,7311.273 293 731 131.355 614 8131.355\,614\,8131.355 614 8 1.102 904 5081.102\,904\,5081.102 904 508 – w/ Init. 8.170 039 698.170\,039\,698.170 039 69 1.273 790 7891.273\,790\,7891.273 790 789 20.861 568 0320.861\,568\,0320.861 568 03 1.214 094 851.214\,094\,851.214 094 85 Ours 707070% 9.180 407 8239.180\,407\,8239.180 407 823 1.272 074 7881.272\,074\,7881.272 074 788 163.676 240 5163.676\,240\,5163.676 240 5 1.061 940 9851.061\,940\,9851.061 940 985 – w/ Init. 8.086 262 248.086\,262\,248.086 262 24 1.273 578 8461.273\,578\,8461.273 578 846 21.151 118 9621.151\,118\,9621.151 118 96 1.213 969 0641.213\,969\,0641.213 969 064 BAF 19.674 893 0519.674\,893\,0519.674 893 05 1.259 772 491.259\,772\,491.259 772 49 125.028 145 1125.028\,145\,1125.028 145 1 1.127 308 9021.127\,308\,9021.127 308 902 – w/ Init. 8.253 078 2028.253\,078\,2028.253 078 202 1.273 636 991.273\,636\,991.273 636 99 19.550 436 619.550\,436\,619.550 436 6 1.213 505 481.213\,505\,481.213 505 48 111 Hz CMax [17] 19.395 190 2319.395\,190\,2319.395 190 23 1.275 866 1481.275\,866\,1481.275 866 148 117.439 798 6117.439\,798\,6117.439 798 6 1.143 852 7451.143\,852\,7451.143 852 745 – w/ Init. 8.254 263 2638.254\,263\,2638.254 263 263 1.289 966 6551.289\,966\,6551.289 966 655 20.627 507 7120.627\,507\,7120.627 507 71 1.223 032 6881.223\,032\,6881.223 032 688 Downsampling 909090% 8.676 346 0658.676\,346\,0658.676 346 065 1.289 459 4391.289\,459\,4391.289 459 439 110.568 366 3110.568\,366\,3110.568 366 3 1.130 495 7731.130\,495\,7731.130 495 773 – w/ Init. 8.184 187 8618.184\,187\,8618.184 187 861 1.289 611 9121.289\,611\,9121.289 611 912 20.620 358 2120.620\,358\,2120.620 358 21 1.223 050 3321.223\,050\,3321.223 050 332 Ours 909090% 8.505 807 8288.505\,807\,8288.505 807 828 1.289 787 0891.289\,787\,0891.289 787 089 87.774 541 8687.774\,541\,8687.774 541 86 1.158 678 1341.158\,678\,1341.158 678 134 – w/ Init. 8.177 036 3548.177\,036\,3548.177 036 354 1.290 007 1851.290\,007\,1851.290 007 185 20.569 191 6120.569\,191\,6120.569 191 61 1.223 052 861.223\,052\,861.223 052 86 BAF 19.568 905 3719.568\,905\,3719.568 905 37 1.275 738 1221.275\,738\,1221.275 738 122 117.554 331 4117.554\,331\,4117.554 331 4 1.143 238 0781.143\,238\,0781.143 238 078 – w/ Init. 8.189 429 3938.189\,429\,3938.189 429 393 1.289 532 0351.289\,532\,0351.289 532 035 20.713 437 520.713\,437\,520.713 437 5 1.223 006 3741.223\,006\,3741.223 006 374
@@ 7.3 Denoising Convergence During Optimization
To further validate the proposed joint estimation approach, we analyze the convergence during the joint estimation using the ECD dataset in Fig. 11 (see also results in Fig. 5). In the first iteration (i.e., initialization), signal and noise events are randomly classified. As optimization proceeds, signal events evolve towards keeping edge structures in the scene, while noise events evolve towards dropping such edge structures (see second and third rows). Also, the IWEs converge to sharp edges with correct motion parameters. This example confirms the efficacy of joint estimation.
The optimization process on an HD-resolution (1280×7201280$\times$ 7201280 × 720 px) real-world dataset, TUM-VIE [33], is shown in Fig. 12. The intermediate progress (d)–(e) demonstrates how motion estimation converges and the denoising performance improves, simultaneously.
@@ 7.5 Full Results on Angular Velocity Estimation
While the quantitative evaluation on angular velocity estimation is summarized in Fig. 6, here, we report the detailed results with different target ratio parameters, also compared with other baselines such as BA Filter [9]. The original CMax degrades due to noise, as reported in previous work (e.g., [3]). The signal-to-noise (S/N) target ratio $\tau$ affects accuracy: the closer it is to the actual value of noise injection, the better the results of the proposed method. The amount of artificial noise injected is around 151515 % for 555 Hz and 333 % for 111 Hz conditions. Although the “true” noise level is unknown due to the original noise in the ECD sequences, our method constantly produces better accuracy and FWL values than the baselines. Please refer to Sec. 4.3 for more discussions about the dependency on initialization and comparison with other baselines. The AUCs for the conditions that we test ($\tau=\{0.9,\ldots,0.7\}$) are 0.700.700.70 (“Ours”) and 0.670.670.67 (“Downsampling”).
@@ 7.6 Quantitative Evaluation of Intensity Reconstruction
Nonetheless, we report non-reference image quality indices for different S/N ratios $\tau$. Figure 13 reports the scores of Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [38] and Naturalness Image Quality Evaluator (NIQE) [39], using Bicycle-ND64-2 sequence (same as Fig. 8). These scores indicate the perceptual quality of images, and smaller is better. Although BRISQUE monotonically increases as the target ratio decreases (i.e., more events are removed), NIQE scores the lowest at $\tau=0.9$, indicating the best quality of the reconstructed image. Although the results potentially suggest that it could estimate the “true” noise ratio in the data using the non-reference indices, which is useful for image reconstruction applications, we leave further evaluation and discussion as future work.
Figure 13: Results of the non-reference image quality indices for image reconstruction. See Fig. 8 for details.